OpenStudy (anonymous):

Assignment is : Write the slope-intercept of the equation of each line. problems are as follows: 6x=5y=-15 I do not understandwhat I am supposed to do; is thier a formula to do this?

OpenStudy (anonymous):

slope=6/5; slope=-a/b if ax+by+c=0